EDUCATION

University of California at Berkeley, Berkeley, CA

PhD in Mathematics expected May, 2000. Advisor: H.W. Lenstra Oberlin College, Oberlin, OH

BA in Mathematics, with high honors, 1990.

RESEARCH EXPERIENCE

Associated Member Technical Staff AT&T Bell Labs, Summers 1990-1995 (except 1992)

TEACHING EXPERIENCE

Graduate Student Instructor

Calculus with Workshop, UC Berkeley, Fall 1998 Calculus Instructor

Summer Bridge Program, UC Berkeley, Summer 1992 Math Instructor

Project Star, UC Berkeley, Summer 1992 Pre-Calculus Teacher

Washington Ethical Society School, Washington, DC, 1985-1986

OTHER PROFESSIONAL EXPERIENCE

Trader/Financial markets researcher Renaissance Technologies Corporation, Stony Brook, NY, 1995-1998

AWARDS AND HONORS

AT&T Cooperative Research Fellowship, Mentor: J.C. Lagarias

1993-1995 National Defense Sciences and Engineering Graduate Fellowship

1990-1993 National Science Foundaion Minority Graduate Fellowship

1990 (declined) John D. Baum Memorial Prize in Mathematics

Oberlin College, 1990 Sigma Xi

Oberlin College, 1990 Phi Beta Kappa

Oberlin College, 1989

PAPERS

“A polyomino tiling problem of Thurston and its configurational entropy”

Math. Monthly, March 1991 “Connected goalies of convex polygons”, (to be submitted)

TALKS “Newton polygons and higher ramification”, February, 1999

Intercity Number Theory Seminar, R.U. Leiden, The Netherlands “Strongly Eisenstein polynomials”, November, 1999

UC Berkeley Number Theory Seminar “A polyomino tiling problem of Thurston”, July, 1991

AT&T Bell Labs Poster session